OpenStudy (anonymous):

A circle of radius 7 is centered at the point (-8, 9). Find a formula for f(q) (Use the variable q for the angle.)

OpenStudy (amistre64):

x^2 + y^2 = r^2 cos^2 + sin^2 = 7^2 49cos(q) + 49sin(q) = 49 ??? hows that

OpenStudy (amistre64):

ack!!... forgot to include the off centerd lol

OpenStudy (amistre64):

i was sooooo close...

OpenStudy (amistre64):

(x+8)^2 +(y-9)^2 = 49 x^2 +16x +64 +y^2 -18y +81 = 49

OpenStudy (amistre64):

cos^2(q) + 16cos(q) + sin^2(q) -18sin(q) +96 = 0 maybe??

OpenStudy (anonymous):

just to parametrize maybe they are looking for x= -8+7cos(q) y= 9+7sin(q) i don't know if that is what is meant maybe?

OpenStudy (amistre64):

hard to tell with the question; I was gonna parametric it, but thought maybe not :)

OpenStudy (anonymous):

Thank you that helped